# Static Detection of Core Structures in Tigress Virtualization-Based Obfuscation Using an LLVM Pass

### v1

@@ 1.1 Research Background
To address these challenges, this paper proposes a method for analyzing VM-obfuscated code using static analysis, which does not require direct code execution. We suggest a method to define and extract core structures at the LLVM (Low-Level Virtual Machine) [9] Intermediate Representation (IR) level. By developing a static analysis-based tool, we aim to accurately identify the core components of a virtual machine—namely the dispatch routine, handler blocks, and the VM region—to provide stable and versatile information for interpreting the original semantics. While this research focuses on the Tigress obfuscator, it is expected to be extensible to other commercial obfuscation tools in the future. The remainder of this paper is organized as follows: Section 2 introduces the research background and related work. Section 3 analyzes the structure of code protected by VM-based obfuscation. Section 4 presents the methodology for identifying core structures based on static analysis. Section 5 verifies the effectiveness of the proposed static analysis through experimental results, and Section 6 concludes the paper and discusses future research directions.
@@ 2.1 VM-based Obfuscation
Figure 1 shows the Control Flow Graph (CFG) of a code sample where VM-based obfuscation has been applied. The blue box at the top represents the block where the dispatcher begins; it can be observed that the switch structure has been transformed into nested conditional statements through compilation. The blocks within the red boxes are handlers, and among them, the purple box in the center is the VM exit handler, which is the block that exits the dispatcher structure.
@@ 2.4 LLVM (Low-Level Virtual Machine)
Figure 2: LLVM Framework structure [9]
@@ 3.1 Dispatcher Structure Analysis
In general, the structure of VM-obfuscated code is described by three key terms: the Virtual Program Counter (VPC), the virtual instruction array, and the dispatcher, which serves as an interpreter to execute them. Among these, the dispatcher acts as the execution framework for the virtual machine based on the VPC, the virtual instruction array, and the handlers. The VPC determines the next handler to be executed, ensuring that the actual instructions corresponding to the obfuscated program are executed in sequence from among the handlers existing in the virtual interpreter array.
@@ 4 Identification of Core Structures Based on Static Analysis
4.1 Contribution of the LLVM IR Analysis Environment
To perform static analysis, an environment based on LLVM 22.0.0 was first established. The analysis was conducted on an Ubuntu 24.04 environment, and the experimental data was generated by applying Tigress 4.0.11 VM-based obfuscation to code written in the C language. In line with the execution of the LLVM Pass, the Tigress obfuscation environment options were set to Linux and Clang. Specifically, original C files were obfuscated using Tigress with specific options, and the resulting obfuscated C code was then converted into LLVM IR using the Clang compiler to perform static analysis on the generated files.
@@ 4.2 Design and Implementation of Feature-Based LLVM Pass
Table 1: Description of Core Roles in VM-based Obfuscated Code
@@
Since understanding the architecture of the virtualized machine is crucial for analyzing VM-obfuscated programs, this paper designs and implements an LLVM Pass that detects the core structures of the obfuscated code through LLVM IR analysis. The core structures defined in this study, as shown in Table 1, include the Dispatch Start, Handler blocks, and the VM Start/End points. This approach enables the static analysis of the control flow of the obfuscated code without direct execution. Consequently, the specific structural features extracted through static analysis were inserted into the LLVM IR in the form of specially designed dummy function calls, allowing these features to be utilized for core structure detection even at the binary code stage [^1] The analysis code developed for this study is available on GitHub(https://github.com/sangjun19/LLVM-Pass/tree/main/VMTag).
@@ 4.2.1 Identification of the Dispatch Start Point
Consequently, this Dispatch S tart point possesses the highest number of exit points (successors). The code in Figure 6 iterates through all basic blocks within function F and designates the basic block with the highest number of successors as the Dispatch Start. Notably, this central hub block appears even in the direct and indirect options of the Tigress tool, despite their threaded code form. This occurs because, since almost identical goto instructions222goto*(_TIG_VZ_gSL9_1_main_CalcRegion_\$pc_inlinse_5[0])); are used at the end of each handler, it was confirmed that the compiler performs a default optimization by repeatedly utilizing a single basic block. In other words, even for direct and indirect options, the Dispatch Start block can be distinguished as the block with the most successors.
@@ 4.2.3 Identification of VM Start/End Points
The VM region represents the entire scope of the code where VM-based obfuscation is actually applied. VM Start is defined as the block that calls or branches to the Dispatch block; we confirmed that this part is indeed the starting point of the VM region by comparing it with the original code in the actual VM-obfuscated code. The VM End point is defined as a handler block that terminates the VM logic of the program instead of branching back to the Dispatch block among all Handler blocks.
5 Experimental Results
@@ 5.1 Experimental Environment and Data
The experiments were conducted on Ubuntu 24.04. To generate virtualization-obfuscated data, we utilized C source codes for a sorting algorithm, factorial, and Fibonacci sequences. The control structures of these codes include nested loops, single loops, and single loops within conditional statements.
We applied three major virtualization options (switch, direct, and indirect) of the Tigress obfuscation tool to these C codes. The obfuscated codes were then compiled using the Clang compiler under various optimization levels: no optimization (-O0) and several optimization options (-O1, -O2, -O3, -Ofast, -Og, -Os, -Oz). Through this process, we generated the corresponding LLVM Intermediate Representation (IR) code.
@@ 5.2 Experimental Results and Discussion
Figure 7 illustrates the results of executing the proposed LLVM Pass on the virtualization-obfuscated code, demonstrating its ability to identify key structures. Table 2 summarizes the detection status of these core structures across the three control structures. In the -O0 environment, our proposed method successfully identified all four core structures within the obfuscated intermediate language: the VM Dispatcher, VM Handlers, and the VM Start and End points.
@@
Unlike previous studies, the proposed method is a technique that utilizes only static analysis, assuming an environment where dynamic information is difficult to obtain. Therefore, we believe that combining it with dynamic analysis would yield even better performance. Furthermore, since the conversion to LLVM IR is essential for performing this static analysis, a lifting stage is required for binary-level analysis. If high-performance and reliable lifters become available in the future, it is expected that the proposed technique can be effectively extended and applied to the field of binary analysis.
@@ 6 Conclusion
This paper researched a method for statically identifying the core structures of VM-obfuscated code. Through this approach, it is possible to confirm the dispatcher and the skeletal structure of the VM by automatically analyzing only the obfuscated code without the need to extract dynamic traces. We utilized an LLVM Pass and analyzed various virtualization options. In the future, we plan to research methods for performing the semantic interpretation of obfuscated code statically, based on these core structure analysis techniques.

### v2

@@ 1.1 Research Background
To address these challenges, this paper proposes a method for analyzing VM-obfuscated code using static analysis, which does not require code execution. We suggest a method to define and extract core structures at the LLVM (Low-Level Virtual Machine) [llvm] Intermediate Representation (IR) level. By developing a static analysis-based tool, we aim to accurately identify the core components of a virtual machine—namely the dispatch routine, handler blocks, and the VM region—to provide stable and versatile information for interpreting the original semantics. While this research focuses on the Tigress obfuscator, it is expected to be extensible to other commercial obfuscation tools in the future. The remainder of this paper is organized as follows: Section 2 introduces the research background and related work. Section 3 analyzes the structure of code protected by VM-based obfuscation. Section 4 presents the methodology for identifying core structures based on static analysis. Section 5 verifies the effectiveness of the proposed static analysis through experimental results, and Section 6 concludes the paper and discusses future work.
@@ 2.1 VM-based Obfuscation
Figure 1 shows the Control Flow Graph (CFG) of a code sample where VM-based obfuscation has been applied. The blue box at the top represents the block where the dispatcher begins; it can be observed that the switch structure has been transformed into nested conditional statements through compilation. The blocks within the red boxes are handlers, and among them, the black box in the center is the VM exit handler, which is the block that exits the dispatcher structure.
@@ 2.4 LLVM (Low-Level Virtual Machine)
Figure 2: LLVM Framework structure
@@ 3.1 Dispatcher Structure Analysis
In general, the structure of VM-obfuscated code is described by three components: the Virtual Program Counter (VPC), the virtual instruction array, and the dispatcher, which serves as an interpreter to execute them. Among these, the dispatcher acts as the execution framework for the virtual machine based on the VPC, the virtual instruction array, and the handlers. The VPC determines the next handler to be executed, ensuring that the actual instructions corresponding to the obfuscated program are executed in sequence from among the handlers existing in the virtual interpreter array.
@@ 4 Identification of Core Structures Based on Static Analysis
4.1 Construction of the LLVM IR Analysis Environment
To perform static analysis, an environment based on LLVM 22.0.0, Ubuntu 24.04. The experimental dataset was generated by applying VM-based obfuscation to C-based source code using the Tigress 4.0.11 obfuscation tool. To isolate the effects of VM-based obfuscation, other obfuscation were excluded. The Tigress environment was configured for Linux and Clang to ensure compatibility with the execution of the LLVM Pass. Specifically, original C files were obfuscated using specific Tigress options, and the resulting obfuscated C code was converted into LLVM IR via the Clang compiler. Static analysis was then performed using theses IR files. While the current analysis focuses on IR derived from source code, it is expected that this approach can be extended to IR obtained through binary code lifting tools in the future.
@@ 4.2 Design and Implementation of Feature-Based LLVM Pass
Table 1: Description of Core Structures in VM-Obfuscated Code
@@
Since understanding the architecture of the virtualized machine is crucial for analyzing VM-obfuscated programs, this paper designs and implements an LLVM Pass that detects the core structures of the obfuscated code through LLVM IR analysis. The core structures defined in this study, as shown in Table 1, include the Dispatch Start, Handler Blocks, and the VM Start/End. This approach enables static analysis of the control flow of the obfuscated code without direct execution. Consequently, the specific structural features extracted through static analysis were inserted into the LLVM IR in the form of specially designed dummy function calls, allowing these features to be utilized for core structure detection even at the binary code stage [^1] The analysis code developed for this study is available on GitHub(https://github.com/sangjun19/LLVM-Pass/tree/main/VMTag).
@@ 4.2.1 Identification of the Dispatch Start Point
Consequently, this Dispatch Start point possesses the highest number of exit points (successors). The pseudocode in Figure 6 iterates through all basic blocks within the function and designates the basic block with the highest number of successors as the Dispatch Start. Notably, this central hub block appears even in the direct and indirect options of the Tigress tool, despite their threaded code form. This occurs because, since almost identical goto instructions222goto*(_TIG_VZ_gSL9_1_main_CalcRegion_\$pc_inlinse_5[0])); are used at the end of each handler, it was confirmed to be due to the compiler’s default behavior of repeatedly utilizing a single common block. Consequently, even in direct and indirect options, the Dispatch Start block can be identified by locating the block with the highest number of outgoing edges.
@@ 4.2.3 Identification of VM Start/End Points
The VM region represents the entire scope of the code where VM-based obfuscation is actually applied. VM Start is defined as the block that calls or branches to the Dispatch Start. We confirmed that this part is indeed the starting point of the VM region by comparing it with the original code in the actual VM-obfuscated code. The VM End point is defined as a handler block that terminates the VM logic of the program instead of branching back to the Dispatch Start among all Handler Blocks.
5 Experiment
@@ 5.1 Experimental Environment and Data
The experimental setup was identical to the previously described LLVM IR analysis environment. The evaluation was conducted using C-based source code featuring various iterative structures, such as sorting algorithms, factorial calculations, and Fibonacci sequences. The control structures of these codes include nested loops, single loops, and single loops within conditional statements.
We applied three major virtualization options (switch, direct, and indirect) of the Tigress obfuscation tool to these C codes. The obfuscated codes were then compiled using the Clang compiler under various optimization levels: no optimization (-O0) and several optimization options (-O1, -O2, -O3, -Ofast, -Og, -Os, -Oz). Through this process, we generated the corresponding LLVM IR code.
@@ 5.2 Experimental Results and Discussion
Figure 7 illustrates the results of executing the proposed LLVM Pass on the VM-obfuscated code, demonstrating its ability to identify key structures. Table 2 summarizes the detection status of these core structures across the three control structures. In the -O0 environment, our proposed method successfully identified all four core structures within the obfuscated intermediate language: the VM Dispatcher, VM Handlers, and the VM Start and End points.
@@
Unlike previous studies, this research distinguishes itself by utilizing static analysis exclusively, without relying on dynamic information. Furthermore, while the current approach targets LLVM IR, it is expected that this technique can be effectively extended to the binary domain once high performance binary to IR liters become available in the future.
@@ 6 Conclusion
This paper researched a method for statically identifying the core structures of VM-obfuscated code. Through this approach, it is possible to confirm the dispatcher and the core structure of the VM by automatically analyzing only the obfuscated code without the need to extract dynamic traces. We utilized an LLVM Pass and analyzed various virtualization options. In the future, based on these core structural analysis methods, we plan to study the code to perform semantic interpretation of obfuscation and to identify related analyses and roles.